### v1

# The role of quantum and classical correlations in shrinking algorithms for optimization

Abstract
The benefit of quantum computing for solving combinatorial optimization problems (COPs) constitutes an open research question. In this work, we study the performance of a shrinking algorithm for COPs. The algorithm leverages correlations extracted from quantum or classical subroutines to recursively simplify the problem. We compare the performance of the algorithm equipped with correlations from the quantum approximate optimization algorithm (QAOA) as well as the classical linear programming (LP) and semi-definite programming (SDP) relaxations. This allows us to benchmark the utility of QAOA correlations against established classical relaxation algorithms. We apply the recursive algorithm to MaxCut  problem instances with up to a hundred vertices at different graph densities. Our results indicate that LP outperforms all other approaches for low-density instances, while SDP excels for high-density problems. Moreover, the shrinking algorithm proves to be a viable alternative to established methods of rounding LP and SDP relaxations. In addition, the recursive shrinking algorithm outperforms its bare counterparts for all three types of correlations, i.e., LP with spanning tree rounding, the Goemans-Williamson algorithm, and conventional QAOA. While the lowest depth QAOA consistently yields worse results than the SDP, our tensor network experiments show that the performance increases significantly for deeper QAOA circuits.
I Introduction
Combinatorial optimization problems (COPs) are highly relevant in diverse sectors across both science and industry. However, they are in general challenging to solve efficiently because of the exponential growth of the solution space with increasing problem size. Traditionally, a variety of classical strategies, including linear programming (LP) [1] and semi-definite programming (SDP) relaxations [2], have been applied to find approximate solutions to COPs. Recently, approaches utilizing quantum resources have emerged as promising alternative solution strategies [3]. Among these, the quantum approximate optimization algorithm (QAOA) [4] is of particular interest because of its universality and adaptive complexity. QAOA is a hybrid quantum-classical variational algorithm designed to find approximate solutions to unconstrained binary optimization problems using noisy-intermediate scale quantum (NISQ) devices and beyond.
It is unclear whether NISQ algorithms can in practice provide speedups over classical methods for solving COPs. In fact, several issues of the near-term variational approaches have been identified, such as barren plateaus [5] and noise of quantum devices [6]. Another issue of QAOA is its locality, i.e., the effect that only qubits that are separated by less than a certain distance in the graph representation of a problem can interact with each other at a given circuit depth. This property can be used to prove limitation results on the performance of QAOA [7, 8, 9, 10].
The recursive QAOA (RQAOA) was introduced by Bravyi et al. [7, 11] to overcome the locality-induced limitations of QAOA. The algorithm operates by recursively simplifying the problem based on correlations between variables obtained from QAOA. The two steps of computing the correlations and fixing the variables are executed iteratively until the problem is fully solved. By introducing new connections between previously unlinked variables, RQAOA overcomes the locality of QAOA. Several works have generalized and extended RQAOA. These proposals include using problem-specific update rules [12, 13], analog quantum devices [12] or a shrinking procedure based on the classical calculation of correlations [14]. It has been shown that those recursive algorithms outperform the original QAOA for many problem instances [15, 11, 16]. However, it is not clear whether the good results are due to the recursive shrinking procedure or the quantum correlations. Here, we aim to give insights in this direction.
This study focuses on evaluating how different classical and quantum methods for calculating correlations affect the performance of a shrinking algorithm similar to RQAOA. The investigated routines for calculating correlations are the classical LP and SDP relaxations and QAOA. Our numerical experiments on instances of MaxCut allow us to benchmark the utility of quantum correlations against correlations obtained from established classical approximation algorithms. Moreover, we find that the shrinking procedure employed here is a viable alternative to traditional rounding routines for LP and SDP relaxations.
The remainder of the paper is organized as follows. In Section II we introduce the MaxCut problem and the different routines (LP, SDP and QAOA) for calculating the correlations. In addition, we present the shrinking algorithm and how it utilizes the calculated correlations. In Section III we analyze the performance of the shrinking algorithm employing different correlation sources, before discussing the implications of our findings in Section IV. In Section V we conclude by suggesting potential future research directions.
II Methods
We begin this section by introducing two mathematical formulations of the MaxCut  problem. Next, we provide details on how the correlations can be computed using the LP, SDP, and QAOA routines. Finally, we present the shrinking algorithm that uses the correlations to recursively simplify the optimization problem.
II-A MaxCut problem and its encoding
Despite the simple structure of the MaxCut problem, it constitutes a popular example of a COP since any quadratic unconstrained binary optimization (QUBO) problem can be transformed into a MaxCut problem [17, 18, 19, 20]. Furthermore, solving it is NP-hard [21], and finding solutions for dense instances with hundreds of variables can already overstrain state-of-the-art algorithms [22].
A MaxCut problem instance is defined by a weighted undirected graph $G=(V,E)$ with vertices $V=\{i\}$, edges $E=\{e\}$ and edge weights $w_{e}$. We denote the number of vertices by $n=|V|$. In MaxCut, we are tasked with finding a subset of nodes that maximizes the weight of edges connecting the chosen node subset and its complement. Formally, we want to find a node partition $W\subseteq V$ such that the edge set $\delta(W)\coloneqq\{ij\in E\mid i\in W,j\in V\setminus W\}$ maximizes its weight defined as $\sum_{e\in\delta(W)}w_{e}$.
The weighted MaxCut problem can also be formulated as maximizing an integer quadratic unconstrained cost function $C(\mathbf{x})$ in the form of
$C(\mathbf{x})=\frac{1}{2}\sum_{ij\in E}w_{ij}(1-x_{i}x_{j}).$ (1)
Here, $\mathbf{x}\in\{-1,1\}^{n}$, and $x_{i}$ indicates whether vertex $i$ is in the subset $W$ or not. Furthermore, $w_{ij}\in\mathbb{R}$ represent the edge weights.
II-B Means of computing correlations
At the core of the algorithm are the correlations between decision variables of a MaxCut problem. Each edge $ij\in E$ in the problem graph is assigned a correlation $b_{ij}\in[-1,1]$. Ideally, we want this correlation to be indicative of the correlation between variables in high-quality solutions of the MaxCut problem. In this case, a large negative (positive) correlation between two variables indicates that in good candidate solutions, these two variables take mostly opposite (equal) values on average. In the language of MaxCut, this translates to an edge predominantly being cut for negative correlations and not cut for positive correlations.
The shrinking algorithm works identically irrespective of the correlations that are used. However, for different situations, different means of computing correlations might be better suited to achieve the best possible results. Thus, in this paper, we compare three means of computing correlations, namely, using LP and SDP relaxations, and QAOA.
II-B1 Linear programming (LP)
For a weighted undirected graph $G=(V,E)$, we define an edge-incidence vector $\mathbf{y}\in\{0,1\}^{|E|}$. Here, 00 corresponds to the edge not being cut, and $1$ represents the edge being cut.
An integer linear programming formulation of the MaxCut problem can be weitten as (cf. [23, 24, 25, 26])
$\begin{split}\max\quad&\sum_{e\in E}\omega_{e}y_{e},\end{split}$ (2a)
$\begin{split}\text{s.t.:}\quad&\sum_{e\in Q}y_{e}-\sum_{e\in C\backslash Q}y_{e}\leq|Q|-1,\\[3.0pt] &\qquad\ \>|Q|\ \text{odd},\ \forall Q\subseteq C\ \text{cycle},\end{split}$ (2b)
$\begin{split}&0\leq y_{e}\leq 1,\ \forall e\in E,\end{split}$ (2c)
$\begin{split}&y_{e}\in\{0,1\},\ \forall e\in E,\end{split}$ (2d)
where equations (2b) are the so-called odd-cycle inequalities, known to be sufficient to define a cut. While the MaxCut  problem defined by Eqs. (2a), (2b), (LABEL:eq:_odd-cycle_d) cannot be solved efficiently in general, the linear relaxation defined by Eqs. (2a)–(LABEL:eq:_odd-cycle_c) can be solved in polynomial time, e.g., by a branch-and-bound approach, and yields an upper bound on the maximum cut [22]. This also means that if the relaxation returns an integer solution, the problem has immediately been solved to optimality by the relaxation.
Analogously to Ref. [14], we compute correlations from an optimum solution $\tilde{\mathbf{y}}\in[0,1]^{|E|}$ to the linear program Eqs. (2a)–(LABEL:eq:_odd-cycle_c) by the means of the affine function
$b_{e}^{\mathrm{LP}}\coloneqq 1-2\tilde{y}_{e}\in[-1,1].$ (3)
The intuition behind these correlations is that we get strong correlations when either the edge weight is large or when the decision to cut an edge does not interfere with other desired cuts. Here, the term interfering stands for the following situation: Say, we are deciding to cut a specific edge. This decision makes it impossible to cut other edges that we might want to cut as well. In such a case we assign a worse correlation to this specific edge to circumvent cutting it.
II-B2 Semi-definite programming (SDP)
The calculation of semi-definite programming correlations is inspired by the seminal approximation algorithm introduced by Goemans and Williamson [2], which is the best known polynomial-time approximation algorithm for MaxCut  problems. The main idea underlying the Goemans-Williamson (GW) algorithm is that we replace the NP-hard integer quadratic problem formulation of MaxCut as introduced in Eq. (1) by a relaxed version of the problem. The relaxation admits a larger solution space that contains all possible solutions to the integer problem. Thus, the optimal solution of the relaxed problem upper bounds the best solution of the integer model by design. Concretely, we replace integer values with multi-dimensional vectors $\mathbf{v}_{i}\in\mathbb{R}^{n}$ with $\norm{\mathbf{v}_{i}}=1$, i.e., to the $(n-1)$-dimensional unit sphere $S_{n-1}$. Using this, we can relax the maximization of Eq. (1) to
$\begin{split}\max_{\{\mathbf{v}_{1},\mathbf{v}_{2},\dots\}}\quad&\frac{1}{2}\sum_{i<j}w_{ij}(1-\mathbf{v}_{i}\cdot\mathbf{v}_{j}),\\ \text{s.t.}\quad&\mathbf{v}_{i}\in S_{n-1}\quad\forall i\in V.\end{split}$ (4)
This model is equivalent to the semi-definite program
$\max_{X}\ \left\{\big{\langle}\frac{1}{4}L,X\big{\rangle}:\text{diag}(X)=\mathbf{{e}},X\succeq 0\right\},$ (5)
where $L\coloneqq\text{diag}(A\mathbf{{e}})-A$ is the Laplacian of the problem graph and $\mathbf{{e}}$ stands for the vector of all ones. Here, $\langle a,b\rangle=\text{tr}(ab^{\text{T}})$ represents the Frobenius inner product of two matrices $a$, $b\in\mathbb{R}^{n\times m}$. This generic semi-definite program can be solved in polynomial time by various solvers such as cvxopt, see, e.g., [27, 28, 29].
Furthermore, $X$ denotes the Gram matrix of $\{\mathbf{v}_{i}\}$. To obtain the vectors $\{\mathbf{v}_{i}\}$, one needs to compute the Cholesky decomposition of the matrix $X$: $X=B^{\text{T}}B$, which has a computational complexity of $O(n^{3})$. In the resulting matrix $B$, the $i$th column corresponds to the vector $\mathbf{v}_{i}$.
We will consider two possibilities for calculating correlations based on SDP relaxations. The first possibility, which we refer to as SDP correlations, is more straightforward and can be directly computed from the vectors. The second option – GW correlations – requires additional evaluations of the MaxCut objective function.
SDP correlations
From the vectors $\{\mathbf{v}_{i}\}$, the correlation for each edge $\{i,j\}\in E$ is obtained by evaluating the dot product
$b_{ij}^{\mathrm{SDP}}=\mathbf{v}_{i}\cdot\mathbf{v}_{j}\in[-1,1],$ (6)
which measures the level of (anti-)alignment of the vectors. This way we assign a large absolute correlation if the vectors are approximately parallelly or anti-parallelly aligned, i.e. the dot product takes a value close to $+1$ or $-1$. Comparing Eqs. (1) and (4) reveals that this relaxed formulation is analogous to the case of integer variables, where an edge is cut if $x_{i}x_{j}=-1$, and not cut for $x_{i}x_{j}=+1$. In contrast to the integer formulation, the relaxed dot product correlations can now take values between $-1$ and $+1$. Furthermore, we can compute these correlations efficiently.
GW correlations
Alternatively, we can obtain correlations by leveraging the hyperplane rounding procedure from the GW algorithm [2]. The rounding procedure maps vectors to an integer solution, yielding an approximation of the maximum cut. This is done by first choosing a random hyperplane, defined by a normal vector $\mathbf{r}\in S_{n-1}$. Next, we assign the vertices $V$ into two partitions $S\subseteq V$ and $\overline{S}=V\,\backslash\,S$ using the criterion:
$\quad S=\{i\ |\ \mathbf{v}_{i}\cdot\mathbf{r}\geq 0\}.$ (7)
As the evaluation of a given candidate cut can be performed efficiently, it is possible to try several random hyperplanes, calculate the resulting cut value, and choose the hyperplane corresponding to the best cut value. It has been shown in Ref. [2] that the cuts obtained this way are guaranteed to achieve an approximation ratio of at least $0.878$.
For a given hyperplane that divides the vectors into two partitions, we define the GW correlations as
$b_{ij}^{\mathrm{GW}}=\begin{cases}\ \frac{1}{2}(\mathbf{v}_{i}\cdot\mathbf{v}_{j}+1)&\text{if}\ i,\ j\ \text{are in the same partition,}\\[4.0pt] \ \frac{1}{2}(\mathbf{v}_{i}\cdot\mathbf{v}_{j}-1)&\text{if}\ i,\ j\ \text{arein opposite partitions.}\end{cases}$ (8)
The GW correlations are designed to combine both the global information about the assignment of the vectors to one of the two partitions defined by the hyperplane, as well as the local information about the (anti-)parallel alignment of the two vectors.
Figure 1: Depiction of the shrinking algorithm used in this work: First, the MaxCut problem instance is modeled according to the input requirement of the chosen routine for computing the correlations. In step a) we use the chosen routine to compute the correlations between variables. Then in b) the problem graph is shrunk to a smaller MaxCut problem by fixing a variable based on the calculated correlations and given update rules. The shrinking is repeated $r$ times before a new set of correlations is calculated for the simplified graph. The steps a) and b) are applied in an alternating manner until the problem is fully simplified. In the final step c) the solution to the original problem is reconstructed based on the fixed variables.
II-B3 QAOA
Finally, we turn to computing the correlations using QAOA. The goal of QAOA is to find an $\mathbf{x}\in\{-1,1\}^{n}$ that maximizes the cost function $C:\{-1,1\}^{n}\rightarrow\mathbb{R}$ of an arbitrary integer quadratic unconstrained optimization problem. Starting from the uniform superposition state $\ket{+}$ defined by
|$\ket{+}=\frac{1}{\sqrt{2^{n}}}\sum_{\mathbf{x}\in\{-1,1\}^{n}}\ket{\mathbf{x}},$ (9)
we apply a sequence of parametrized unitaries to prepare the state
|$\ket{\Psi(\boldsymbol{\beta},\boldsymbol{\gamma})}=e^{-i\beta_{p}H_{M}}e^{-i\gamma_{p}H_{C}}\cdots e^{-i\beta_{1}H_{M}}e^{-i\gamma_{1}H_{C}}\ket{+}.$ (10)
Here, $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})$, $\boldsymbol{\gamma}=(\gamma_{1},\dots,\gamma_{p})$ are real-valued parameters and $p\in\mathbb{N}$ is the depth. $H_{C}$ denotes the cost Hamiltonian that encodes the optimization problem as
$H_{C}\ket{\mathbf{x}}=C(\mathbf{x})\ket{\mathbf{x}}\quad\forall\mathbf{x}\in\{-1,1\}^{n},$ (11)
and $H_{M}$ is the so-called mixer Hamiltonian
$H_{M}=\sum_{i=1}^{n}X_{i},$ (12)
where $X_{i}$ is the Pauli-X operator applied on qubit $i$. We then use a classical subroutine that optimizes the parameters ($\boldsymbol{\beta}$, $\boldsymbol{\gamma}$) such that the expectation value
$F(\boldsymbol{\beta},\boldsymbol{\gamma})=\bra{\Psi(\boldsymbol{\beta},\boldsymbol{\gamma})}H_{C}\ket{\Psi(\boldsymbol{\beta},\boldsymbol{\gamma})}$ (13)
is maximized. Thus, we hope to prepare a superposition of high quality candidate solutions.
As follows from Eq. (1), for MaxCut the cost Hamiltonian $H_{C}$ can be written as
$H_{C}=\frac{1}{2}\sum_{ij\in E}w_{ij}(I-Z_{i}Z_{j}),$ (14)
where $I$ is the identity operator and $Z_{i}$ is the Pauli-Z operator acting on qubit $i$. After maximizing the expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$, correlations $b_{ij}$ between nodes $i$ and $j$ connected by an edge can be defined as the expectation value with respect to the QAOA state:
$b_{ij}^{\mathrm{QAOA}}=\langle Z_{i}Z_{j}\rangle\equiv\bra{\Psi(\boldsymbol{\beta_{\text{opt}}},\boldsymbol{\gamma_{\text{opt}}})}Z_{i}Z_{j}\ket{\Psi(\boldsymbol{\beta_{\text{opt}}},\boldsymbol{\gamma_{\text{opt}}})},$ (15)
where $\boldsymbol{\beta_{\text{opt}}}$ and $\boldsymbol{\gamma_{\text{opt}}}$ stand for the optimized parameters after maximization.
In our experiments we simulated QAOA for MaxCut problems with depths $p\in\{1,2,3\}$. For depth $p=1$ the QAOA correlations $b_{ij}^{\mathrm{QAOA}}$ and thus also the expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$ can be expressed analytically as shown in Ref. [30]. In the present work, at $p=1$ we maximize the QAOA expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$ by first applying a rough grid search over the possible parameter space of $\boldsymbol{\beta},\boldsymbol{\gamma}$ before using the gradient-based BFGS algorithm [31] with the best parameters from the grid-search as initial parameters.
For depths $p>1$ we use the Qtensor tensor network library [32, 33, 34] to compute the expectation values and correlations. The basis of the simulation is mapping the quantum circuits that need to be evaluated to tensor networks. This can be done by interpreting a quantum state of $n$ qubits as tensor from $(\mathbb{C}^{2})^{\otimes n}$ and a quantum gate as a tensor with input and output indices for each qubit it acts on [35]. Here, an input index corresponds to the output index of the previous gate. A contraction of the resulting tensor network leads to an exact simulation of the quantum circuit. Because the required computational effort for calculating the result of such a network depends strongly on the order in which the indices are contracted over, it is crucial to optimize the order of contraction [36, 37]. In this work, we use a method of finding a good contraction order based on a line graph representation of the tensor network as introduced in Refs. [33, 32].
Then, we use gradient ascent to find the optimal variational parameters $\boldsymbol{\beta_{\text{opt}}},\boldsymbol{\gamma_{\text{opt}}}$ that maximize $F(\boldsymbol{\beta},\boldsymbol{\gamma})$. To ensure the reliability of the optimization, it is essential to start gradient ascent from suitable initial parameters $\boldsymbol{\beta}_{\text{i}},\boldsymbol{\gamma}_{\text{i}}$ [38]. In this work, we used the so-called fixed-angle conjectures as initial parameters for the tensor network simulations guaranteeing a good performance [39]. Although the conjectures are derived for $k$-regular graphs, we also use them as initial parameters for non-regular graphs with geometries similar to $k$-regular problems. These graphs arise from applying the shrinking steps (introduced in the next subsection) to $k$-regular graphs.
II-C Shrinking procedure
MaxCut is particularly suitable for shrinking algorithms, as there is a natural way of reducing the problem such that the shrunk problem is still a valid MaxCut problem. However, extensions of the shrinking presented here can be applied to other problems as well [12].
The shrinking procedure used in this work is similar to RQAOA [7, 11] as well as the algorithm introduced in Ref. [14]. For a better overview, a schematic of the algorithm is shown in Fig. 1. In general, the problem can be divided into three major steps a) to c). We will give a brief high-level overview before explaining all the steps in detail in the following paragraphs. After modeling the MaxCut problem, we first start with calculating correlations between the decision variables of the problem in step a). Then in step b), we use these correlations to successively reduce the number of nodes in the problem graph. We do this by combining two nodes, generating a new MaxCut instance that has one node fewer. The reduced problem is equivalent to the original problem under the additional constraint implied by the correlation. Importantly, since the shrunk problem is again a valid MaxCut problem, we can calculate new correlations for the shrunk problem using the same method as for the original problem. This can be done either after every shrinking step or after a given number $r$ of shrinking steps. Lastly, in step c), we recreate a solution to the original instance from a solution to the shrunk instance by undoing the shrinking steps. In the following, the three steps are described in more detail.
Computing the correlations
At the core of our scheme is a method for computing the correlations between the decision variables of a given MaxCut problem instance. These are obtained by the routines presented in Section II-B. We want to emphasize that the shrinking procedure is independent of how the correlations are calculated. However, for different means of computing correlations, the algorithm’s performance will vary. This is because the correlations are calculated from approximations and will hence typically not be optimal. However, when using perfect correlations for every shrinking step, i.e., correlations obtained from an optimal cut of the problem, the shrinking algorithm is guaranteed to return the optimal cut.
Shrinking
Next, we present how correlations are used to guide the shrinking procedure. After computing the correlations, we find the correlation with the largest absolute value, where ties are broken randomly. This way, the algorithm first selects the edges that have the strongest tendency to be cut or not cut.
Specifically, once the maximal correlation, say $b_{ij}$, is determined, the algorithm identifies both the vertex pair $i,j$ and the sign $\sigma_{ij}$ of the correlation:
$\sigma_{ij}\coloneqq\text{sign}(b_{ij}).$ (16)
Then, for positive (negative) $\sigma_{ij}$, these two vertices are fixed to lie in the same (opposite) partition, reducing the total number of vertices by one (see panel b) in Fig. 1). To ensure consistency, the weights of the graph are updated such that the new graph $G^{\prime}=(V^{\prime},E^{\prime})$ is consistent with the original graph $(V,E)$ together with the newly imposed constraint for the nodes $i$ and $j$. For a node $i\in V$, its neighborhood is defined as $\mathcal{N}(i)\coloneqq\{k\in V\ |\ ik\in E\}$. For all neighbors $k\in\mathcal{N}(i)$ of node $i$, the new weights $\omega_{jk}^{\prime}$ are defined by
$w_{jk}^{\prime}=\begin{cases}\ \omega_{jk}+\sigma_{ij}\omega_{ik},&\quad\text{if}\ jk\ \in E\\[4.0pt] \ \sigma_{ij}\omega_{ik},&\quad\text{if}\ jk\ \notin E,\end{cases}$ (17)
while all other weights remain unaltered. In comparison to the original graph $(G,E)$, the updated graph $(G^{\prime},E^{\prime})$ lacks node $i$, and in some cases, two edges were merged into a single one:
$\begin{gathered}V^{\prime}=V\,\backslash\,\{i\}\\ E^{\prime}=\left(E\cup\{jk:\ k\in\mathcal{N}(i)\}\right)\,\backslash\,\left\{ik:\ k\in\mathcal{N}(i)\right\}.\end{gathered}$ (18)
An example of these update rules is given in Fig. 1 b). It is important to keep track of the association of vertices in the shrunk graph with those of the original graph, in order to be able to reconstruct the final solution at the end.
On a more technical note, it is possible that after several shrinking steps, a node $u$ belonging to the considered correlation $b_{su}$ does not exist anymore because it had already been combined with another node $v$. This situation is depicted in panel b) of Fig. 1. There, we first calculate the correlations $b_{uv},b_{su},b_{ut},b_{tv}$ for the input graph on the left. The correlations are already sorted in descending order of absolute values. Then we shrink the graph on the left side of box b) according to the chosen correlation $b_{uv}$. This step yields the output graph on the right-hand side, where the node $u$ has been merged onto node $v$. The next best correlation is now $b_{su}$. However, node $u$ does not exist in the current graph. In this case, we simply interpret the correlation $b_{su}$ as the correlation $b_{sv}$. Note that we must take care of the sign used in the previous step to ensure consistency. To this end, we set $b_{sv}=\sigma_{uv}\cdot b_{su}$. Furthermore, analogous rules apply if both nodes $i$ and $j$ have already been combined with other nodes. For the special case that the two to-be shrunk nodes $i$ and $j$ already belong to the same node $k$, the shrinking step is skipped, and the algorithm continues by using the next best correlation. This skipping does not count as step in the procedure.
Since the shrinking step does not fix individual variables to be in a certain partition, but rather fixes the relation between two variables, the shrunk problem remains a valid MaxCut problem. This allows us to calculate correlations of the shrunk problem in the same manner as for the original graph. Intuitively, it is likely that correlations computed using the shrunk graph provide a better source of information for the shrinking step, as they have been specifically tailored to the shrunk graph. Therefore, it makes sense to introduce recalculations of the correlations for the shrunk graph after a given interval of $r$ shrinking steps, as also shown in Fig. 1 b). With these newly obtained correlations, we shrink the reduced problem for another $r$ steps until the correlations are recalculated again. For correlations obtained by QAOA and a recalculation interval of $r=1$, the shrinking algorithm corresponds to the RQAOA algorithm from Refs. [7, 11]. Note that recalculation can be performed using any of the means for computing the correlations discussed here.
In this work, the shrinking procedure is applied until the problem graph has been shrunk to a size of two nodes, when it becomes trivial to solve. A potential extension of our algorithm performs the shrinking procedure until the problem graph has been reduced to a sufficiently small size such that an exact classical solver can find an optimal solution.
Reconstruction of the solution
Finally, from a solution of the shrunk graph together with the history of all shrinking steps, a solution to the original graph is reconstructed. Starting from a solution of the shrunk graph, one simply has to backtrack through the performed shrinking steps. At each shrinking step, we either add the shrunk node to the vertex partition or exclude it, depending on the sign of the shrinking.
III Results
Here, we present numerical experiments of applying the proposed shrinking algorithm using the different methods of computing correlations introduced in Sec. II. We first introduce the MaxCut  problem instances and performance metrics used in this work. Afterward, we compare the performance of the shrinking algorithm with the different means of calculating correlations to their corresponding bare approximation algorithms. Next, we investigate the influence of recalculating the correlations by studying the algorithm’s performance at different recalculation intervals. Finally, we employ tensor network methods to analyze the benefits of using deeper QAOA circuits to generate improved quantum correlations.
III-A Problem instances and approximation ratio for benchmarks
For the experiments in this study, two families of graphs are used. Most experiments are performed on Erdős-Rényi random graphs with a hundred nodes and different densities. An Erdős-Rényi graph for a given number of nodes and density $d$ is an undirected and unweighted graph with random edges. Each pair of nodes is connected by an edge with probability $d$.
Since simulating deep QAOA circuits requires computational resources that scale prohibitively with the graph density, we use sparse graphs for the tensor network experiments. Specifically, we use random 3-regular graphs with $50$ nodes. This type of graph has been thoroughly studied before in the context of QAOA for MaxCut  problems [39, 40, 41, 42] and has been used in hardware experiments because of their modest requirements on the qubit connectivity of quantum devices [43, 44].
Throughout the results section we use the approximation ratio as an indicator for the performance of the shrinking algorithm. We define the approximation ratio as
$R_{A}\coloneqq\frac{S_{A}}{S_{G}}.$ (19)
Here, $S_{A}$ is the cut size obtained by the investigated algorithm. $S_{G}$ stands for the cut size retrieved from Gurobi [45], a state-of-the-art commercial solver. For each instance, we run Gurobi on a single core (Dual AMD Rome 7742) and terminate the optimization when the value of the optimization objective has not changed in an hour. We emphasize that in this way, it is not guaranteed that the Gurobi optimizer solves the problem instances to optimality, especially in the case of dense Erdős-Rényi graphs. However, we believe that the Gurobi benchmark is sufficiently strong for the purposes of this study.
III-B Comparison to the bare underlying algorithm
Figure 2: Median approximation ratio $R_{A}$ of the bare LP, GW and QAOA ($p=1$) algorithms as well as their shrinking counterparts with recalculation intervals $r=1$ and $r=\infty$. The algorithms are applied to 80 different randomly generated Erdős-Rényi graphs of size 100 for each of the densities $0.1$, $0.4$ and $0.8$. The lower and the upper error bars represent the first and third quartiles, respectively.
[…]
We begin our analysis of the results by comparing the shrinking algorithm using various correlations to the approximate solutions obtained by the bare underlying algorithms used to compute the correlations. We first explain how the performance of the bare underlying algorithms is evaluated.
For the bare LP algorithm, we calculate the correlations for each edge in the manner described in Section II-B1 and then round these correlations to an integer solution by using a maximum spanning tree rounding heuristic. This is a well-established heuristic, previously used in Refs. [23, 46, 47]. For SDP, we use the standard Goemans-Williamson algorithm as the bare algorithm. Here, we do the hyperplane rounding (as described in Section II-B2) $15$ times and return the best solution. For QAOA, the graphs are too large to perform a classical simulation of the standard quantum algorithm. Efficient sampling from the output state of a QAOA $p=1$ circuit with one hundred qubits already exceeds the capabilities of classical hardware [48]. To this end, we calculate the expectation value of the energy returned by $p=1$ QAOA using analytical formulae from [30].
[…]
Starting with LP, we observe that the shrinking with no recalculations outperforms the spanning tree heuristic for all of the considered graph densities, even though the correlations are identical in both cases. In addition, there is a clear improvement in performance when increasing the amount of recalculations performed (and, thus, decreasing $r$). This also holds true for SDP and QAOA. However, in contrast to SDP and QAOA, which show a roughly constant approximation ratio for all densities, LP performs significantly better at lower densities than on high densities.
The bare GW algorithm and the SDP shrinking algorithm perform well across all densities. The shrinking algorithm with recalculation interval $r=1$ performs best with median approximation ratios above $99\%$ for all densities. However, the bare GW algorithm slightly edges out the shrinking algorithm with no recalculations.
Finally, for QAOA the shrinking algorithm with recalculation interval $r=1$ (equivalent to RQAOA) sees a significant improvement over the bare algorithm: For the different densities, the median approximation ratio increases roughly from around $90\%$ for the bare QAOA to approximately $99\%$ for the $r=1$ shrinking algorithm. This behavior is in line with previous results from the literature [7, 11]. Interestingly, QAOA shows the largest relative improvement compared to the bare algorithm of the three different means of computing correlations. The bare algorithm and the shrinking with no recalculations perform quite similarly. However, the error bars indicate that the quality of the shrinking with no recalculations varies significantly more than the quality of the standard QAOA algorithm.
III-C Effect of the recalculation interval and problem density
[…]
First, let us focus on the effect of the recalculation interval in general before looking at the impact of the problem density on the performance. For all three means of computing correlations, there is a clear trend that more recalculations improve the solutions. However, in the case of LP correlations, for low densities, the $r=10$ shrinking algorithm slightly outperforms its $r=1$ variant. While the performance differences for LP and the SDP between $r=10$ and $r=1$ are small, QAOA shows a significant increase for $r=1$. For all, SDP, LP and QAOA, the recalculation intervals $50$ and $\infty$ result in far worse approximation ratios than smaller recalculation intervals. This highlights the importance of recalculating correlations.
Turning now to the effect of problem density, the most striking dependence is observed for LP correlations. There, the approximation ratio for low-density instances is significantly better than for any other type of correlation. This is also in correspondence with the general good performance of the bare LP algorithm for sparse graphs, as also shown in other publications like Ref. [22].
Another notable feature in Fig. 3 is the systematic increase of the median approximation ratio for densities above $0.3$. This increase is likely an artifact of a worse quality of the reference solutions obtained via Gurobi, rather than improved performance of the shrinking algorithm. While for sparser instances, Gurobi can certify that the returned solution is optimal, the same is not true for denser instances. This leads us to believe that our choice of the benchmark is causing the increasing trend of the approximation ratio observed across all correlation types. However, it is important to note the present setup still allows us to perform comparisons between different correlation routines.
Thus, we see that the performance of the LP shrinking algorithm decreases very fast when the density increases from $0.05$. In contrast, the SDP and QAOA correlations lead to a more consistent performance of the shrinking algorithm across the entire range of densities. In addition, the SDP correlations clearly outperform $p=1$ QAOA for most densities, while the latter, in turn, outshines LP correlations for densities higher than $0.2$.
III-D Performance of the GW correlations
After comparing SDP, LP and QAOA as means of computing correlations, we evaluate the GW correlations separately to ensure a fair comparison: While the previously analyzed strategies simply compute correlations using the problem graph, the algorithm for the GW correlations from paragraph II-B2 evaluates the MaxCut objective function several times directly while searching for a good hyperplane. Thus, it utilizes information not accessible to the other routines for computing correlations. Nonetheless, these correlations can still be computed efficiently in polynomial time.
[…]
Furthermore, the GW correlations for $r=50$ have a similar performance as the SDP correlations for $r=1$. From this, we conclude that the best approximation ratio for the SDP correlations can also be achieved by the GW correlations with less computational resources. Finally, it is worth mentioning that GW correlations with a recalculation interval of $r=10$ yield a better median approximation ratio than with $r=1$ for all densities. The reason for this is that the GW correlations use additional information, which has a larger positive impact on the performance when correlations are used for several shrinking steps without recalculations. This finding is particularly interesting since less computational resources yield a better result, in contrast to the analysis for the SDP correlations in the previous chapter. There, the performance steadily increases with the number of recalculations. As discussed more in detail in Section IV, this behaviour could result from the additional information about the cut value to which only the GW correlations have access.
[…]
III-E Better quantum correlations lead to improved performance
Next, we analyze the performance of the proposed shrinking algorithm informed by improved quantum correlations. This is achieved by considering QAOA circuits at depths $p\in\{2,3\}$. The quality of QAOA is known to improve with increasing $p$ [4].
[…]
Fig. 5 showcases the results of our experiments, where the approximation ratio is computed with respect to the Gurobi solutions as defined in Eq. (19). We note that in this case Gurobi was able to certify the optimality of the solutions. For comparison, we also plot the results obtained via LP and SDP correlations. In the figure the heights of the bars correspond to the median approximation ratio across all problem instances and runs of an algorithm. The upper and lower error bars represent the third and first quartiles of the approximation ratios obtained, respectively. We run the shrinking algorithm multiple times for each instance due to the random tie-breaking when choosing a correlation in a shrinking step. This can lead to different final solutions. This effect is particularly important to QAOA, where at low depths $p$ the correlations depend only on the local neighborhood of a given edge. For 3-regular graphs, there is a limited number of distinct local neighborhoods, which leads to many ties in the correlations [30, 49].
Crucially, results in Fig. 5 confirm the intuition that deeper QAOA circuits should generate better correlations. This is evident from the increased performance of the bare algorithms as the depth $p$ increases. Moreover, this steady performance increase transfers over to the performance of the shrinking algorithm, irrespective of the amount of recalculations performed. Furthermore, we again observe that the performance of the shrinking algorithm improves when recalculating the correlations after every shrinking step. In fact, we are able to optimally solve all of the considered graphs already at $p=2$ if the recalculation of correlations is performed after each shrinking step. Nevertheless, it is reasonable to expect that for larger instances higher depths would be required. Finally, we note that the shrinking algorithm with no recalculations ($r=\infty$) on average outperforms the bare algorithm, even at higher depths.
Because of the low density of the 3-regular graphs, the problem instances are in the density regime where the shrinking algorithm of LP and its classical counterpart have the best performance of the different algorithms, as we have seen in section III-C. Thus, all variants of LP reach the optimal solutions. Furthermore, the size of the problems is small enough that also all three SDP algorithms attain very good approximation ratios such that the median approximation ratios are at $100\%$. However, the SDP shrinking variant with $r=\infty$ solves more problem instances perfectly than the bare algorithm. Only the shrinking version with $r=1$ is able to achieve optimal results for all MaxCut instances.
[…]
IV Discussion
In this work, we have extended and analyzed algorithms that solve a combinatorial optimization problem by recursively shrinking it. We compared the utility of different methods for computing correlations, both quantum and classical, and thus provided classical benchmarks that quantum-informed shrinking will need to surpass in order to be useful in practice. We remark that our proposed shrinking procedure can be regarded as a standalone heuristic that improves the performance of the bare underlying algorithm. We ran numerical simulations for the shrinking procedure applied to the MaxCut problem.
Interestingly, we observe that the shrinking algorithm not only increases the performance of the quantum algorithm QAOA, as already shown in [7, 11], but can also significantly improve upon the performance of classical algorithms like the Goemans-Williamson algorithm or the linear programming relaxation of MaxCut. Furthermore, the proposed shrinking procedure can serve as a rounding heuristic for the relaxed LP or SDP. Similarly to the spanning tree heuristic, it rounds perfect correlations to the optimal solution. However, our studies indicate that it delivers better integer solutions for fractional correlations than spanning tree rounding.
Applying the shrinking algorithm to Erdős-Rényi graphs of various densities shows that the shrinking algorithm employing LP correlations perform extraordinarily well for low densities, whereas the performance quickly plummets as the density increases. Hence, this suggests a very simple decision metric about the utility of the LP correlations: For solving low-density instances, the LP correlations should be used, but for higher-density instances, other correlations are expected to achieve better results. This sensitivity of the LP to the problem density can be explained by the definition of the relaxation model. Whereas the other relaxations use node-assignment as the decision variables, LP uses decision variables based on edges. This means that the number of variables for a constant number of nodes in the LP increases significantly when increasing the density of the problem graph.
The SDP and the QAOA correlations have a more constant performance where the approximation ratio slightly increases with increasing density. As mentioned before, this can be caused both by an increase in the quality of the shrinking algorithm solution or a decrease in the performance of the reference Gurobi solution.
Furthermore, for all three methods of computing correlations, we observe that more recalculations tend to yield better results. However, this improved performance comes at the cost of an increase in the computational resources required to run the algorithm. Notably, the performance of the LP and SDP-informed shrinking algorithm only improves slightly when recalculating correlations in every step $(r=1)$ compared to every ten steps $(r=10)$.
The correlations from the GW-rounding perform best over the entire range of problem instances considered in this work. Interestingly, while there is still a trend that the recalculations improve the performance, the best results with GW correlations are returned for a recalculation interval of $r=10$. In contrast, for the SDP correlations the performance increases monotonously with the number of recalculations, even though in both cases the correlations are based on the same algorithm. Intuitively, the peak in performance at an intermediate value of $r$ can be understood in terms of the competing effects of recalculations of correlations and GW rounding. Our results generally indicate that increasing the number of recalculations improves the performance. However, performing a recalculation of correlations also requires a new choice of the rounding hyperplane. This new choice of the hyperplane might cause inconsistencies with previously performed shrinking steps. We believe that this trade-off between the amount of recalculations and new choices of hyperplanes is responsible for the observed best performance at $r=10$.
Furthermore, we have confirmed that better quantum correlations improve the performance of the shrinking algorithm by simulating higher-depth QAOA for the correlations used in the algorithm. This result indicates that further improvement of quantum hardware will enable performance gains for shrinking algorithms, thereby providing clues for whether quantum algorithms can provide practically relevant methods for combinatorial optimization.
V Outlook
This study explores the role and practicality of classical and quantum correlations for a recursive shrinking algorithm applied to MaxCut  instances. A natural extension of our work would be to consider other optimization problems. The focus of our paper is primarily the comparison of different sources correlations and their potential use. However, there are several avenues towards increasing the algorithm’s performance.
As discussed in [14], one way to improve the performance of the shrinking algorithm is to not shrink completely until it becomes trivially solvable. Instead, the algorithm should shrink the problem to a size where it can be solved to optimality by a different exact solver. By doing so, we can avoid mistakes during the last few shrinking steps, leading to an improved solution. Moreover, one could envision devising better update rules or using backtracking techniques to rectify shrinking steps that caused a decrease in performance [12].
Furthermore, an interesting property of the graph shrinking algorithm is that with each shrinking step both the graph connectivity and its edge weights change. This is particularly interesting when considering persistency checks for MaxCut, as it is sometimes possible to make assumptions about whether an edge is cut or not by simply considering the MaxCut  graph itself [26, 50]. As the graph changes with each shrinking step, it is possible to conduct these kinds of checks at every shrinking step, even before calculating the correlations, and making a decision for the next step with certainty based on these checks. This can also only increase the performance of the algorithm, but, of course, the checks need to be executed efficiently. On a high-level, this can be seen as an example of a problem-specific shrinking step, that leverages the specifics of the optimization problem [12].
Another way of increasing the performance is the possibility of updating the correlations when shrinking steps without recalculation are performed. For example, if two edges are combined during a shrinking step, the new correlation assigned to this edge could be taken to be e.g., the mean of the correlations belonging to the original edges.
Furthermore, there are numerous other means of obtaining correlations beyond the ones discussed here. For instance, much interest in quantum approaches to combinatorial optimization has been devoted to analog devices [51, 52]. Previously, shrinking algorithms using correlations from an analog device have been proposed in Ref. [12]. We believe that a systematic study of the performance of shrinking algorithms using quantum correlations from analog and gate-based quantum hardware compared to classical benchmarks could provide a useful stepping stone towards a better understanding of the utility of quantum computers for solving COPs.
Acknowledgments
FW thanks Frauke Liers for many fruitful and encouraging discussions. JRF acknowledges Libor Caha and Aron Kerschbaumer for insightful discussions. VF and MP thank Johannes Frank for inspiring dialogues. JRF and MP thank the BMW team for their support. VF, FW, LP and CM are supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), project QuaST.